Suppose the Isoquant and Isocost curves of a manufacturing company are given as Q=4K^0.5 L^0.5 and 4K + 8L= 320 respectively, where r is the price of capital and w refers to the price of labour. How many units of L and K the firm should employ to maximize profit?
1
Expert's answer
2021-09-22T07:28:15-0400

Solution:

The profit-maximizing rule is the price of input resource must be equal to its marginal revenue product.

Q = 4K0.5L0.5


"\frac{MP_{L} }{w} = \frac{MP_{K} }{r} = \frac{MP_{L} }{MP_{K}} = \frac{w }{r}"


MPL = "\frac{\partial L} {\partial Q}" = 2K0.5L-0.5


MPK = "\frac{\partial K} {\partial Q}" = 2K-0.5 L0.5


"\frac{MP_{L} }{MP_{K}} = \frac{w }{r}"


= "\frac{2K^{0.5} L^{0.5}}{2K^{-0.5} L^{0.5}} = \frac{8 }{4}"


="\frac{K }{L} = \frac{8 }{4}"

K = 2L

Plug into isocost line:

320 = 4K + 8L

320 = 4(2L) + 8L

320 = 8L + 8L

320 = 16L

L = 20

K = 2L = 2"\times"20 = 40

 

The units of L and K the firm should employ to maximize profit are = 20 and 40